# How Teachers Can Use Large Language Models and Bloom’s Taxonomy to Create Educational Quizzes

### v1

@@ 1 Introduction
[…]
@@ 2 Background
Generate a question from the following passage: <<<>>>
@@
Generate a multiple-choice question from the following passage: <<<>>>
@@ 3.1 Contexts
Candidate educational questions were generated from a set of $24$ passages sourced from Wikipedia, with each passage containing $5$ context paragraphs, $6$ to $9$ sentences in length. The total number of input contexts used to generate questions is, thus, $120$. The length of these contexts was determined by empirical results in preliminary work. This set of passages was manually gathered. The specific Wikipedia articles were generally chosen via hyperlinks from the domain’s main Wikipedia or glossary page in order to ensure they were relevant to the foundations of the given domain. Two domains were used: biology (BIO) and machine learning (ML). More than one domain is included in order to take steps towards demonstrating domain-agnostic results. Each domain has $12$ passages, totaling $60$ input contexts. The contexts underwent minor pre-processing before use: this included removal of citations, hyperlinks, footnotes and phonetic spellings, re-formatting of full sentence bullet-point lists into paragraphs, and other minor data cleaning steps.111 The input contexts, more details about Bloom’s taxonomy, the human authored few-shot examples, all of the generated candidates and quizzes, the annotator demographics, and more can be found at https://anonymous.4open.science/r/EQG˙in˙practice-2752/README.md.
@@ 3.2 Simple Prompting Strategy
[…]
@@ 3.3 Controlled Prompting Strategy
[…]
@@
[…]
@@ 4.2 Quiz Writing Experiments
[…]
@@
[…]

### v2

@@ 1 Introduction
((b)) Controlled prompting strategy generation examples.1 All 6 levels of Bloom’s taxonomy have generations for each input context, omitted here to save space.
@@ 2 Background
Generate a question from the following passage: <<>>
@@
Generate a multiple-choice question from the following passage: <<>>
@@ 3.1 Contexts
Candidate educational questions were generated from a set of $24$ passages sourced from Wikipedia, with each passage containing $5$ context paragraphs, $6$ to $9$ sentences in length. The total number of input contexts used to generate questions is, thus, $120$. The length of these contexts was determined by empirical results in preliminary work. This set of passages was manually gathered. The specific Wikipedia articles were generally chosen via hyperlinks from the domain’s main Wikipedia or glossary page in order to ensure they were relevant to the foundations of the given domain. Two domains were used: biology (BIO) and machine learning (ML). More than one domain is included in order to take steps towards demonstrating domain-agnostic results. Each domain has $12$ passages, totaling $60$ input contexts. The contexts underwent minor pre-processing before use: this included removal of citations, hyperlinks, footnotes and phonetic spellings, re-formatting of full sentence bullet-point lists into paragraphs, and other minor data cleaning steps. [^1] The input contexts, more details about Bloom’s taxonomy, the human authored few-shot examples, all of the generated candidates and quizzes, the annotator demographics, and more can be found at https://anonymous.4open.science/r/EQG˙in˙practice-2752/README.md.
@@ 3.2 Simple Prompting Strategy
((b)) Simple prompting strategy generation examples.1 There are actually 6 generations, omitted here to save space.
@@ 3.3 Controlled Prompting Strategy
The controlled prompting strategy uses a pedagogical question taxonomy to generate questions with different learning goals in mind. Bloom’s taxonomy is a popular framework for categorizing learning objectives in educational materials (Krathwohl 2002). The taxonomy contains six levels of learning, arranged in a hierarchical order from ‘lower’- to ‘higher’-level thinking skills: remembering, understanding, applying, analyzing, evaluating, and creating.1 Bloom’s taxonomy helps teachers design instructional content that targets specific learning goals.
@@
The controlled prompting strategy also uses few-shot learning. Following success with preliminary experimentation, five-shot learning (i.e., five examples within the prompt) is used. As seen in Figure 1, a single example contains a prompt and six questions, one for each level of Bloom’s taxonomy. The examples were handcrafted by a domain expert, and reviewed by another expert for both question quality and adherence to the intended taxonomic level. Each domain had five sets of contexts and examples, totaling $10$ contexts and $60$ questions.1
@@ 4.2 Quiz Writing Experiments
There were $24$ quiz writing teachers, $12$ in each domain. The BIO teachers were found through the freelance platform Upwork and have at least a high-school level of teaching/tutoring experience. The ML teachers were recruited through word-of-mouth at the institutions of the first author and have at least a university level of teaching/tutoring experience. All teachers were fairly monetarily compensated and signed a consent form before participating. They are all proficient in the English language, and are from relatively diverse demographics.1 The teachers completed a training module where they were provided with example quizzes in their domain and were guided through the creation of each quiz type to ensure they understood the task at hand.
@@
+ Simple Generations+ Controlled GenerationsInput Passage AInput Passage BInput Passage CTeacherHandwritten QuizSimple QuizControlled Quiz